Circular No. 6209

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X-RAY NOVA 1994 IN SCORPIUS S. Sazonov and R. Sunyaev, on behalf of the GRANAT/WATCH Team,

report: “GRO J1655–40 is very bright, according to GRANAT/WATCH in the spectral band 8–20 keV. After the bright outburst of late July–early Aug. (IAUC 6196, 6201, 6205), the source became relatively weak (250 ± 100 mCrab) by Aug. 8. Since that date, its flux gradually increased to 800 ± 200 mCrab on Aug. 16 and to 1400 ± 100 mCrab on Aug. 19.”

S. N. Zhang, B. A. Harmon, W. S. Paciesas, and G. J. Fishman, Marshall Space Flight Center, NASA, report for the Compton Observatory BATSE Team: “The x-ray/radio transient source GRO J1655–40 has been undergoing another bright outburst in hard x-rays, immediately after the last outburst (IAUC 6196) reached a minimum flux level of about 60 mCrab (20–100 keV) on Aug. 9 (IAUC 6205). The energy spectrum can be well fitted by a power law with significant flux detectable up to 200 keV. The peak flux of the current outburst was 600 ± 70 mCrab (20–100 keV) on Aug. 17. The flux has since declined to ∼ 250 ± 50 mCrab on Aug. 19. The photon spectral index changed from −2.1 on Aug. 10 to −2.5 on Aug. 17 and to −1.8 on Aug. 19, consistent with the spectral evolution seen previously from this source. We note that the 8- to 20-keV flux detected by WATCH/GRANAT continued to rise since Aug. 17 (see above). The current activity from July 22 appeared in phase with the previous three major outbursts with peak flux above 1 Crab, which were separated by 120 days.”

GRS 1915+105 Sazanov and Sunyaev also write: “The outburst of GRS 1915+105

detected by GRANAT/WATCH on 19 July (IAUC 6196) is still continuing as of Aug. 19. During the outburst, the source’s flux at 8–20 keV has been ∼ 700 mCrab with no significant variations in intensity detected.”

Zhang et al. also report: “Hard x-ray activity from GRS 1915+105 has continued since a previous report (IAUC 6204). The flux in the band 20– 100 keV peaked on Aug. 11 at 220 ± 25 mCrab, following the intense radio flare on Aug. 10 (IAUC 6204), then decreased rapidly until the source was undetectable on Aug. 14. We note that the 8- to 20-keV flux (see above) continued to increase after Aug. 11, suggesting a hard-to-soft transition in the spectrum.”

1995 August 21 Daniel W. E. Green